# EoRA: Fine-tuning-free Compensation for Compressed LLM with Eigenspace Low-Rank Approximation

### v4

@@ 6 Conclusion
In this work, we present EoRA, a novel fine-tuning-free approach that rapidly boosts the task-specific accuracy of compressed LLMs using minimal calibration data, while offering greater flexibility by relaxing compression format constraints. By projecting compression errors into the task-specific eigenspace of activations, EoRA uses eigenvalues to guide SVD, aligning approximation error with layer-wise compression loss—without any gradient-based training. EoRA achieves strong results across language, commonsense, and math reasoning tasks, outperforming prior low-rank methods Yao et al. (2024); Yuan et al. (2023); Liao et al. (2024). Its training-free design allows quick adaptation to various accuracy-latency trade-offs, and it remains robust under quantization, reducing memory overhead. Additionally, it can serve as a strong initialization for LoRA fine-tuning. Overall, EoRA is a scalable, efficient solution for improving compressed LLMs across diverse deployment settings, with potential extensions to new architectures and modalities.
@@ A.1 Sparsity Error Compensation
EoRA 6.75 37.54 27.53 10.91
@@ A.2 Quantization Error Compensation
EoRA 5.75 39.50 27.20 15.08
A.3 Sparse & Quantization Error Compensation
@@
EoRA 6.89 35.06 27.06 9.86
Here, we examine the feasibility of applying EoRA to compensate for ultra-compressed models that undergo both pruning and quantization. Specifically, we prune LLaMA2-7B/13B and LLaMA3-8B to 2:4 sparsity and quantize them to 4-bit. We set the ranks of both EoRA and SVD to 128 to compensate for the pruning and quantization errors. Similarly to our previous findings, LLaMA3-8B is the least resilient to compression, experiencing a significant drop in both perplexity for language generation and accuracy on commonsense and math reasoning tasks. Notably, the accuracy on ARC-C plummets to $18.33\%$ and MathQA to $19.89\%,$ which is worse than random guessing. However, compensating for the sparsity and quantization errors with EoRA significantly improves the accuracy of these compressed models, reducing perplexity by up to $73.55$ and boosting accuracy by $12.88\%$/$9.60\%$/$10.16\%$ on ARC-C/MathQA/GSM8K tasks. Additionally, EoRA consistently outperforms both ZeroQuant-V2 and Act-S across LLaMA2 and LLaMA3. For instance, EoRA exceeds ZeroQuant-V2 in compensating the compressed LLaMA2-13B on ARC-C by $1.79\%$ and on MathQA by $1.77\%,$ narrowing the accuracy gap with the uncompressed model to just $2.85\%$ on MathQA. Overall, we find that EoRA tends to offer greater accuracy recovery when addressing more aggressive compression settings, ensuring the plausibility of adopting EoRA for mitigating severe compression error.
@@ A.4 Compatibility With Various Compression Methods
EoRA 14.04 34.81 30.05 11.52
In this section studies the generalizability and compatibility of EoRA with different pruning methods beyond SparseGPT. We adopt Wanda [Sun et al., 2024], a method that prunes weights with the smallest magnitudes scaled by their corresponding input activations. For these compression methods, we adhere to the calibration set construction detailed in 4.1, and maintain the same settings when utilizing EoRA to address compression errors. We evaluate EoRAon LLaMA3-8B pruned with Wanda to 2:4 structured sparsity. The ranks of all low-rank compensation methods are set to 128. Table 7 demonstrates that EoRAconsistently outperforms every training-free methods, both ZeroQuant-V2 and Act-S, in improving accuracy across all the tasks. For example, EoRA achieves accuracy gains of 7.77%/4.96%/10.76% on ARC-C/MathQA/GSM8K which is 5.04%/3.32%/10.01% over the improvement brought by Act-s. Furthermore, EoRA outperforms ApiQ on both ARC-C and MathQA by 2.9% and 0.44%. Overall, these findings underscore the effectiveness and generalizability of EoRA across different compression techniques.
A.5 Compensation With Different Ranks
@@
EoRA 40.61 29.17 17.51
A.6 Influence of Different Calibration sizes on EoRA
Model Compression Format #Calib MathQA $\uparrow$
LLaMA3-8B W4 16 36.62
32 36.93
64 37.21
W3 16 26.33
32 27.57
64 29.11
2:4 16 30.05
32 30.18
64 29.91
A.7 Fine-tuning Compressed Models with EoRA
Table 10: Finetune the compressed LLaMA3-8B models with various compression settings and different initialization of the low-rank matrices for Commonsense/Math reasoning tasks.
@@ A.7 Fine-tuning Compressed Models with EoRA
LLaMA3-8B Full-precision - w/o finetuning 50.42 40.10
@@
EoRA 48.54 54.67
GPTQ W4 w/o finetuning 45.90 34.07
@@
EoRA 55.46 56.04
GPTQ W3 w/o finetuning 20.90 22.37
@@
EoRA 47.44 53.90
A.8 Inference Speed Evaluation
Table 11: Comparison of the average per-token latency (batch size 1) for 128-token generation on LLaMA3-70B between full-precision and GPTQ + EoRA (rank=128) with and without our custom EoRA kernel.
Format EoRA Kernel Latency Speedup
FP-16 - 60ms 1x
3-bit - 35ms 1.7x
No 54ms 1.1x
Yes 43ms 1.4x
4-bit - 38ms 1.6x
No 61ms 0.9x
Yes 51ms 1.2x
In language generation, the model produces tokens sequentially, making matrix-vector multiplications the primary factor impacting the inference latency. Consequently, we build our custom EoRA kernel on top of GPTQ’s low-bit quantized matrix vector product kernel, pre-allocating the shared output prior to matrix vector multiplication and integrating the full-precision matrix vector multiplication of $B$ into the quantized kernel reducing redundant memory access. We show the inference speedup of our proposed EoRA kernel in the above Table.11.

### v5

@@ 6 Conclusion
In this work, we present EoRA, a novel fine-tuning-free approach that rapidly boosts the task-specific accuracy of compressed LLMs using minimal calibration data, while offering greater flexibility by relaxing compression format constraints. By projecting compression errors into the task-specific eigenspace of activations, EoRA uses eigenvalues to guide SVD, aligning approximation error with layer-wise compression loss—without any gradient-based training. EoRA achieves strong results across language, commonsense, and math reasoning tasks, outperforming prior low-rank methods (yao2024zeroquantv2; yuan2023asvd; liao2024apiq). Its fine-tuning-free design allows quick adaptation to various accuracy-latency trade-offs, and it remains robust under quantization, reducing memory overhead. Additionally, it can serve as a strong initialization for LoRA fine-tuning. Overall, EoRA is a scalable, efficient solution for improving compressed LLMs across diverse deployment settings, with potential extensions to new architectures and modalities.
@@ A.1 Sparsity Error Compensation
EoRA (Ours) 6.75 37.54 27.53 10.91
Table 4 reports more detailed comparison for sparsity error compensation.
@@ A.2 Quantization Error Compensation
EoRA (Ours) 5.75 39.50 27.20 15.08
Table 5 reports more detailed comparison for quantization error compensation. It is worth noting that EoRA-enhanced models outperform smaller models quantized at higher precision, while high-precision models usually outperform low-precision models. For example on GSM8K (Table 5), 3-bit LLaMA2-13B (4.62) performs worse than 4-bit LLaMA2-7B (9.93), but EoRA-enhanced 3-bit LLaMA2-13B (15.08) outperforms 4-bit LLaMA2-7B. This highlights EoRA’s advantage in terms of greater flexibility: rather than shrinking the model architecture, it enables more effective compression of larger models while preserving higher accuracy.
A.3 Sparsity & Quantization Error Compensation
@@
EoRA (Ours) 6.89 35.06 27.06 9.86
Here, we examine the feasibility of applying EoRA to compensate for ultra-compressed models that undergo both pruning and quantization, as shown in Table 6. Specifically, we prune LLaMA2-7B/13B and LLaMA3-8B to 2:4 sparsity and quantize them to 4-bit. We set the ranks of both EoRA and SVD to 128 to compensate for the pruning and quantization errors. Similarly to our previous findings, LLaMA3-8B is the least resilient to compression, experiencing a significant drop in both perplexity for language generation and accuracy on commonsense and math reasoning tasks. Notably, the accuracy on ARC-C plummets to $18.33\%$ and MathQA to $19.89\%,$ which is worse than random guessing. However, compensating for the sparsity and quantization errors with EoRA significantly improves the accuracy of these compressed models, reducing perplexity by up to $73.55$ and boosting accuracy by $12.88\%$$9.60\%$$10.16\%$ on ARC-C/MathQA/GSM8K tasks. Additionally, EoRA consistently outperforms both ZeroQuant-V2 and Act-S across LLaMA2 and LLaMA3. For instance, EoRA exceeds ZeroQuant-V2 in compensating the compressed LLaMA2-13B on ARC-C by $1.79\%$ and on MathQA by $1.77\%,$ narrowing the accuracy gap with the uncompressed model to just $2.85\%$ on MathQA. Overall, we find that EoRA tends to offer greater accuracy recovery when addressing more aggressive compression settings, ensuring the plausibility of adopting EoRA for mitigating severe compression error.
@@ A.4 Compatibility With Various Compression Methods
EoRA (Ours) 14.04 34.81 30.05 11.52
In this section, we study the generalizability and compatibility of EoRA with different pruning methods beyond SparseGPT. We adopt Wanda (sun2023simple), a method that prunes weights with the smallest magnitudes scaled by their corresponding input activations. For these compression methods, we adhere to the calibration set construction detailed in 4.1, and maintain the same settings when utilizing EoRA to address compression errors. We evaluate EoRAon LLaMA3-8B pruned with Wanda to 2:4 structured sparsity. The ranks of all low-rank compensation methods are set to 128. Table 7 demonstrates that EoRAconsistently outperforms every fine-tuning-free method, both ZeroQuant-V2 and Act-S, in improving accuracy across all the tasks. For example, EoRA achieves accuracy gains of 7.77%/4.96%/10.76% on ARC-C/MathQA/GSM8K which is 5.04%/3.32%/10.01% over the improvement brought by Act-s. Furthermore, EoRA outperforms ApiQ on both ARC-C and MathQA by 2.9% and 0.44%. Overall, these findings underscore the effectiveness and generalizability of EoRA across different compression techniques.
A.5 Influence of Different Calibration sizes on EoRA
Model Quantization Format #Calib. Calib. Time (mins) MathQA $\uparrow$
LLaMA3-8B FP16 - - 40.10
W4 16 6.40 36.62
32 7.04 36.93
64 8.03 37.21
128 10.40 37.46
256 14.43 37.60
512 21.11 37.30
W3 16 6.33 26.33
32 7.20 27.57
64 8.16 29.11
128 11.33 30.34
256 14.17 30.21
512 20.89 30.40
We conducted ablation studies to assess how different calibration set sizes—{16, 32, 64, 128, 256, 512}—affect EoRA’s performance in compensating for 4-bit and 3-bit quantized LLaMA3-8B models on MathQA. As shown in Table 8, increasing the number of calibration samples from 16 to 256 yields a moderate improvement in accuracy (from 36.62 to 37.60) for the W4 model. However, further increasing the calibration size to 512 leads to a slight decline, indicating that the accuracy gain saturates beyond a certain threshold. A similar trend is observed for the W3 model, where performance improves steadily up to 128 samples, after which the benefit plateaus. These findings suggest that while EoRA can leverage additional calibration data to improve accuracy, its performance remains stable and robust even with limited calibration.
A.6 Inference Speed Evaluation
Table 9: Comparison of the average per-token latency (batch size 1) for 128-token generation on LLaMA3-70B between full-precision and GPTQ + EoRA with and without our custom EoRA kernel.
Format EoRA Rank EoRA Kernel Latency $\downarrow$ Speedup $\uparrow$
FP-16 - - 60ms 1x
3-bit - - 35ms 1.7x
64 No 52ms 1.2x
64 Yes 44ms 1.4x
128 No 54ms 1.1x
128 Yes 43ms 1.4x
256 No 58ms 1x
256 Yes 48ms 1.3x
4-bit - - 38ms 1.6x
64 No 60ms 1x
64 Yes 49ms 1.2x
128 No 61ms 1x
128 Yes 51ms 1.2x
256 No 63ms 1x
256 Yes 53ms 1.1x
In language generation, the model produces tokens sequentially, making matrix-vector multiplications the primary factor impacting the inference latency. Consequently, we build our custom EoRA kernel on top of GPTQ’s low-bit quantized matrix vector product kernel, pre-allocating the shared output prior to matrix vector multiplication and integrating the full-precision matrix vector multiplication of $B$ into the quantized kernel reducing redundant memory access. We show the inference speedup of our proposed EoRA kernel in Table 9. The first row shows the FP16 latency (60ms), followed by the 3-bit quantized-only model (35ms). The remaining rows present EoRA latencies at different ranks, both with and without our custom CUDA kernel, alongside 3-bit and 4-bit quantization. As shown in the table, our fused EoRA kernel significantly improves inference speed for both 3-bit and 4-bit quantized models. While there remains some overhead compared to the quantized-only baseline, even with our kernel, using no kernel results in significantly higher latency—sometimes exceeding FP16—due to activation movement overhead, as discussed in Section 4.5. These results underscore the practical deployability of EoRA when paired with our optimized kernel.
A.7 Compensation With Different Ranks
@@
EoRA (Ours) 40.61 29.17 17.51
Table 10 reports more detailed comparison for error compensation with different ranks.
A.8 Fine-tuning Compressed Models with EoRA
Table 11: Fine-tune the compressed LLaMA3-8B models with various compression settings and different initialization of the low-rank matrices for Commonsense/Math reasoning tasks.
@@ A.8 Fine-tuning Compressed Models with EoRA
LLaMA3-8B Full-precision - w/o fine-tuning 50.42 40.10
@@
EoRA (Ours) 48.54 54.67
GPTQ W4 w/o fine-tuning 45.90 34.07
@@
EoRA (Ours) 55.46 56.04
GPTQ W3 w/o fine-tuning 20.90 22.37
@@
EoRA (Ours) 47.44 53.90
Table 11 reports more detailed comparison for fine-tuning compressed models.
A.8.1 Ablation: Fine-tuning with different numbers of training data
Table 12: Ablation study on the effect of using different proportions of the dataset for fine-tuning 2:4 pruned LLaMA3-8B models with varying low-rank matrix initializations on Commonsense/Math reasoning tasks.
Model Dataset Ratio LoRA initialization ARC-C $\uparrow$ MathQA $\uparrow$
LLaMA3-8B - - 50.42 40.10
100% QLoRA 41.30 45.42
LoftQ 43.68 48.77
EoRA (Ours) 48.54 54.67
50% QLoRA 38.56 40.23
LoftQ 41.46 42.51
EoRA (Ours) 46.41 48.91
30% QLoRA 36.77 36.71
LoftQ 39.76 40.60
EoRA (Ours) 43.85 44.79
In this section, we show that fine-tuning with the EoRA-compensated model is robust to various ratios of training data. We follow the setting in Sec. 4.4 on compressed LLaMA3-8B models with 2:4 sparsity compression. As shown in Table 12, using EoRA for initialization consistently outperforms both standard and SVD initialization across various dataset ratios, with accuracy improvements (ARC-E/ARC-C/MathQA) of $3.24\%$$4.95\%$$6.4\%$ and $1.85\%$$4.09\%$$4.19\%$ over SVD when fine-tuning using $50\%$ and $30\%$ training data, respectively.
@@ A.9 Quantizing EoRA To Further Reduce Memory Overhead
Table 13 reports more detailed comparison for quantizing EoRA. For example, when a 512-rank EoRA is quantized from 16-bits to 4-bit on 2:4 pruned LLaMA3-8B, the accuracy drops are only $0.43\%$ on ARC-C while the total model size reduces by $16.49\%$ (11.70 GB $\rightarrow$ 9.77 GB). Additionally, compared to the original uncompensated 2:4 pruned model, quantizing EoRA of rank 128/512 improves accuracy by $4.4\%$$11.4\%$ with a total model size increase of just $2\%$ (9.12 GB $\rightarrow$ 9.28 GB) / $7\%$ (9.12 GB $\rightarrow$ 9.77 GB). For 3-bit quantized LLaMA3-8B compensated with a 4-bit quantized EoRA of rank 128/512 achieves $10.6\%$$19.1\%$ accuracy improvements, with a corresponding model size increase of only $3\%$ (4.63 GB $\rightarrow$ 4.78 GB) / $14\%$ (4.63 GB $\rightarrow$ 5.28 GB). Please see Section 4.5 for more analyses.
A.10 EoRA on More Tasks
Table 14: Comparison of 4-bit and 3-bit quantized LLaMA3-8B on the LLM summarization task.
Model Quantization Format Compensation Method CNN/DailyMail (ROUGE-Lsum) $\uparrow$
LLaMA3-8B 4-bit - 0.1672
ZeroQuant-V2 0.1798
Act-S 0.1786
ApiQ 0.1804
EoRA (Ours) 0.1812
3-bit - 0.0650
ZeroQuant-V2 0.0970
Act-S 0.1286
ApiQ 0.1357
EoRA (Ours) 0.1463
LLM Summarization.
We evaluate EoRA and baseline methods on restoring the summarization capability of quantized models using the testset of CNN/DailyMail — a widely used English-language corpus containing over 300k news articles authored by CNN and Daily Mail journalists. This dataset supports both extractive and abstractive summarization, and we adopt ROUGE-Lsum as the evaluation metric, where higher values indicate better summary quality. We set the rank to 128 for all the methods. For Act-S, ApiQ, and EoRA, we use 128 calibration sentences from WikiText2. The results are shown in Table 14. Notably, EoRA consistently outperforms all baselines across both 4-bit and 3-bit quantized LLaMA3-8B models. Specifically, it improves the ROUGE-Lsum score from 0.1672 to 0.1812 in the 4-bit setting, and from 0.0650 to 0.1463 in the 3-bit setting—demonstrating substantial recovery of summarization performance. These results highlight that EoRA remains effective and competitive in practical, real-world scenarios such as summarization.
Table 15: Comparison of 4-bit and 3-bit quantized LLaMA3.2-3B on multi-task language understanding.
Model Quantization Format EoRA Rank MMLU $\uparrow$
LLaMA3.2-3B FP16 - 54.19
4-bit - 24.16
32 52.53
64 52.49
128 52.93
3-bit - 22.89
32 39.08
64 38.83
128 39.68
Language Understanding.
We tested EoRA on 4-bit quantized LLaMA-3.2-3B, varying the rank from 8 to 128. Using 128 WikiText2 calibration samples, we observe substantial recovery—even rank 8 lifts accuracy from 24.16 to 52.20, as shown in Table 15. This result highlights both the effectiveness of EoRA in compensating for quantization errors in smaller-scale models and its robustness across a range of low-rank settings.
A.11 GPTQ: Channel-wise vs. Group-wise Quantization
Table 16: Comparison of 4-bit and 3-bit group-wise quantized LLaMA3-8B. The group size is set as 128. EoRA rank is set as 128.
Model Quantization Format Compensation Method MathQA $\uparrow$
LLaMA3-8B FP16 - 40.10
W4-groupsize 128 - 38.34
ZeroQuant-V2 38.92
Act-S 38.49
ApiQ 38.90
EoRA (Ours) 39.16
W3-groupsize 128 - 32.52
ZeroQuant-V2 32.39
Act-S 33.33
ApiQ 34.80
EoRA (Ours) 35.10
We initially adopt channel-wise quantization for GPTQ to remain consistent with the original GPTQ setup. More importantly, our goal is to showcase the robustness of EoRA in recovering from even severe quantization errors. That said, we conduct additional experiments using group-wise quantization (group size = 128) to evaluate EoRA’s effectiveness in this more commonly used setting. Table 16 reports results on MathQA for 4-bit and 3-bit group-wise quantized LLaMA3-8B models, referred to as W4-groupsize 128 and W3-groupsize 128, respectively. We compare EoRA against prior compensation methods and set the rank to 128. These results highlight EoRA’s consistent advantage over existing methods in compensating for group-wise quantization. In particular, for the W4-group size 128 configuration, EoRA is able to recover nearly all of the lost accuracy, achieving performance comparable to the original full-precision model.
A.12 Layer-wise Discrepancy Analysis
Table 17: Comparison of layer-wise discrepancy (the less the better) on $q$ projector of LLaMA2-7B.
Method Layer 0 Layer 5 Layer 10 Layer 15 Layer 20 Layer 25 Layer 30
GPTQ (4-bit) 2.3 11.3 13.4 12.6 10.7 10.9 11.5
ZeroQuant-V2 1.7 10.2 10.8 11.3 9.7 9.2 9.9
Act-S 2.0 9.7 9.4 10.3 8.4 7.7 8.9
ApiQ 1.5 7.8 8.2 7.4 6.2 5.9 5.1
EoRA (Ours) 1.4 8.5 8.1 7.2 5.8 6.0 5.4
We follow the layer-wise discrepancy analysis setting of the ApiQ (liao2024apiq) and run the analysis on the $q$ projector of LLaMA2-7B. We compare the discrepancy of layers [0, 5, 10, 15, 20, 25, 30] of different compensation methods of rank 128 and show that EoRA effectively reduces layer-wise discrepancy compared to existing baselines, as shown in Table 17. Notably, EoRA maintains consistently low output activation errors across all layers, especially in the later ones (e.g., 5.8 at layer 20 and 5.4 at layer 30), showing significant improvements over ZeroQuant-V2 and Act-S. EoRA also matches or outperforms ApiQ, demonstrating its effectiveness. Importantly, EoRA accomplishes this with dramatically less compute overhead—typically completing in just a few minutes—whereas ApiQ often requires several hours of optimization. This highlights EoRA’s practical advantage as a highly efficient solution for layer-wise error minimization for recovering the error of low-bit quantization.
A.13 More Comparisons with Recent Low-Rank Approaches
Table 18: Comparison of 4-bit and 3-bit quantized LLaMA3-8B for more recent low-rank approaches.
Model Quantization Format Compensation Method MathQA $\uparrow$
LLaMA3-8B - - 40.10
4-bit - 34.07
FWSVD 35.64
ZeroQuant-V2 36.51
Act-S 35.84
ApiQ 36.18
LQER 35.46
LRC 36.40
CALDERA 36.70
QERA 35.90
SLiM 35.90
OATS 36.01
EoRA (Ours) 37.21
3-bit - 22.37
FWSVD 26.30
ZeroQuant-V2 26.43
Act-S 25.42
ApiQ 26.86
LQER 25.60
LRC 28.64
CALDERA 28.10
QERA 25.32
SLiM 25.91
OATS 25.30
EoRA (Ours) 29.11
Recent Low-Rank approaches can be broadly categorized into three groups: 1) activation-statistics-based scaling methods, including SLiM (mozaffari2024slim), OATS (zhang2024oats), LQER (zhang2024lqer), and QERA (zhang2024qera), 2) iterative low-rank compensation approaches, which are LRC (scetbon2024low) and CALDERA (saha2024compressing), and 3) training-based methods, including LR-QAT (bondarenko2024low) and SLoPe (mozaffari2024slope).
The first group—SLiM, OATS, LQER, and QERA—scales the compression error based on activation statistics prior to applying SVD for low-rank approximation. The underlying intuition is that input channels with higher magnitudes (i.e., outliers) are considered more important, and as a result, their corresponding weight entries should be prioritized over those linked to lower-magnitude activations. All four methods differ slightly in how the scaling diagonal matrix is constructed, but they are fundamentally similar to the ASVD-based scaling baseline already discussed in our paper (see Section 4.1). For example, LQER (zhang2024lqer) mitigates quantization error by scaling the residual using activation statistics—specifically, the maximum average magnitude per input channel—before applying SVD. Another example, SLiM (mozaffari2024slim), introduces a low-rank compensation method that also leverages activation statistics (specifically, average absolute activation values) to scale residuals. Although SLIM is proposed alongside a custom quantization scheme, we isolate and evaluate its saliency-based low-rank adapter strategy. Although these methods vary in how they scale the compression error, none of them are guaranteed to minimize the layer-wise compression error directly—an issue we discussed in Section 2 of our paper. In contrast, EoRA is explicitly formulated to minimize this objective (as detailed in Section 3).
The second group, including iterative low-rank compensation approaches, mainly lack the flexibility as EoRA. For example, LRC (scetbon2024low) requires iterative updates to weights and low-rank modules, leading to task-specific quantized models. In contrast, EoRA only adapts the low-rank modules, allowing a shared compressed backbone and easier integration with multi-adapter frameworks (multilora). Although LRC offers a closed-form solution, it assumes is full-rank—an assumption that often fails and requires extra modification steps that may introduce noise and instability. EoRA avoids this by only requiring to be symmetric, improving numerical stability and robustness. Another example, CALDERA (saha2024compressing), employs an iterative optimization strategy that updates both the quantized weights and the low-rank matrices, using a closed-form solution. However, because it requires modifying the quantized weights during this process, CALDERA is less efficient for multi-task scenarios, where separate quantized models would need to be maintained for each task. EoRA, on the other hand, avoids this limitation by making only the low-rank components task-specific, while keeping the quantized backbone fixed and shared across tasks. This decoupled design allows for easy integration with existing multi-adapter inference frameworks (multilora) and significantly improves the practicality of EoRA for real-world applications.
The third group requires gradient-based training. For example, LR-QAT (bondarenko2024low) is a quantization-aware training (QAT) method. While it also uses low-rank modules, it operates in the fine-tuning regime, combining cross-entropy loss with final-layer output alignment (akin to knowledge distillation). SLoPe (mozaffari2024slope) primarily targets improving pre-training efficiency and enhancing the accuracy of compressed models through training. As such, it aligns more closely with quantization-aware training (QAT) as well. In contrast, EoRA is a post-training quantization (PTQ) method that requires no gradient updates to the quantized model. In the LLM compression community (frantar2022gptq; lin2024awq), it is standard practice to distinguish between PTQ and QAT approaches, as they serve different purposes and are not typically benchmarked against each other.
We summarize the empirical comparison on 4-bit and 3-bit quantized LLaMA3-8B models evaluated on MathQA in Table 18. All methods use rank of 128 and identical calibration data (see Section 4.1). To fairly compare CALDERA with EoRA, we adapt the CALDERA method by fixing the quantized weights and only updating the low-rank matrices. This degenerates the iterative process into a two-step approximation: first approximate the down-projection matrix, followed by the up-projection. Once the quantized weights are fixed, additional iterations do not change the approximation.
As the results show, EoRA consistently outperforms other low-rank compensation methods, particularly those based on activation statistics like LQER, SLiM, etc. This highlights the advantage of EoRA’s mathematical property, which directly minimizes the layer-wise compression error rather than relying on heuristics derived from activation statistics. While LRC offers better accuracy than heuristic scaling approaches, it still lags behind EoRA due to its reliance on less stable approximations. These findings highlight EoRA’s practicality and effectiveness for post-training quantized LLMs. From the results, we observe that CALDERA performs well—outperforming both Act-S and ApiQ—but EoRA still consistently achieves higher accuracy. We attribute this to EoRA’s single-step optimization that jointly solves for both projection matrices, whereas CALDERA’s sequential two-step process may accumulate slightly more approximation error.
Overall, EoRA consistently surpasses all the recent low-rank methods, including both activation-statistics-based approaches and iterative low-rank compensation approaches. Its effectiveness stems from its ability to directly minimize layer-wise compression error, eliminate the need for heuristic magnitude-based scaling, and utilize a single-step optimization process. These strengths underscore both the theoretical robustness and practical efficiency of EoRA compared to existing methods.
A.14 EoRA on 2-bit Quantization
Table 19: Comparison of 2-bit GPTQ-quantized LLaMA3-8B on MathQA.
Model Quantization Format Fine-tuning Strategy MathQA $\uparrow$
LLaMA3-8B 2-bit (GPTQ) - 18.22
LoftQ 35.80
EoRA 36.89
RILQ 37.60
RILQ + EoRA 38.90
We also evaluate a more challenging setting, 2-bit quantization, where RILQ (lee2025rilq) is one of the state-of-the-art error compensation methods. RILQ adopts standard backpropagation (i.e., fine-tuning), utilizing a combination of cross-entropy loss and final-layer output alignment (similar to knowledge distillation). As outlined in the RILQ paper, RILQ uses gradient descent to collectively tune all adapters, minimizing the discrepancy between full-precision and quantized activation outputs of the final layer. In addition, as quoted from the RILQ paper, RILQ also incorporates a causal language modeling objective with Ground Truth in the optimization of low-rank adapters. Therefore, it is not appropriate to directly compare EoRA to RILQ as methods in the same category. However, since EoRA can act as an effective initialization for subsequent fine-tuning (as detailed in Section 4.4), it is complementary to RILQ rather than competing with it. To examine this synergy, we performed experiments where EoRA was first used for initialization—following the setup outlined in Section 4.4—and then applied RILQ’s fine-tuning objective to a 2-bit GPTQ-quantized LLaMA3-8B model on the MathQA dataset. The results are shown in Table 19. As expected, RILQ outperforms LoftQ in compensating for quantization error in 2-bit models. Moreover, when initialized with EoRA, further fine-tuning with RILQ’s objective yields an additional 1.3% improvement over RILQ alone.